# Modeling the Solar System as an Observed Multi-Transit System I: Characterization Limits from Analytic Timing Variations

### v1

@@ Abstract
Planetary systems with multiple transiting planets are beneficial for understanding planet occurrence rates and system architectures. Although we have yet to find a solar system (SS) analog, future surveys may detect multiple terrestrial planets transiting a Sun-like star. In this work, we simulate transit timing observations of our Solar System as viewed from a distance and based on the actual orbital motions of Venus and the Earth-Moon Barycenter, as influenced by the other SS bodies, with varying noise levels and observing durations. We then retrieve the system’s dynamical parameters using an approximate N-body model for transit timing shifts while considering four possible plane-parallel configurations: two planets, three planets, four planets, and five planets.
We demonstrate that – with the retrieval applied to simulated transit times of Venus and EMB – we can: 1) detect Jupiter at high significance (up to 90-s timing noise); 2) measure the masses and orbits of both transiting planets (mass-ratios are down to 4-8% uncertainty for the 3-planet model) ; 3) detect Mars with more than $5\sigma$ given very high level precision (10s of seconds). To accurately characterize Jupiter, we require timing precisions of better than 30 seconds and survey durations longer than 22 years. Accurate retrieval of Mars is possible when the baseline is longer than 25 years. Additionally, while Jupiter’s mass is underestimated in most of our simulated cases, the addition of Mars improves the posterior mass, suggesting that unseen terrestrials could interfere in the characterization of multi-planetary systems if they are nearly resonant to transiting planets.
Ultimately, these simulations will help to guide future missions – such as PLATO, Nautilus, LUVOIR, and Ariel – in detecting and characterizing exoplanet systems analogous to our Solar System.
keywords:
astrobiology – planets and satellites: detection, planets and satellites: general
pubyear: 2023pagerange: Modeling the Solar System as an Observed Multi-Transit System I: Characterization Limits from Analytic Timing VariationsA.2
@@ 3 Results
Figure 5: Upper panel: Likelihood profiles for Mars-like (orange curve) and Jupiter-like (firebrick curve) planets fit to the transit times of two terrestrial planets. The times used here were generated over 30 yrs with 30-sec noise. Lower panels: Comparison of likelihood (line) and posterior probability (histogram) of orbital periods near the measured values for Mars (left) and Jupiter (right). We report the median systematic uncertainty, $\sigma_{sys}$, from the posterior sampling of each configuration, indicating that the four-planet model gives a better fit.
@@ 3.4 Effect of timing noise and observing baseline
For the $\mathcal{H}_{PPP}$ case with 30-year duration, we plot how timing precision affects our posterior masses in the top panel of Figure 6. Although an increase in injected noise leads to larger mass uncertainties for a given model, the measured masses for planets b and c remain consistent across all of the models. The lower panel shows how timing precision affects the masses in the $\mathcal{H}_{PPPP}$ case, where available. When the error is $\approx 20$ seconds, the Interestingly, planet b’s mass
When we examine our entire simulation set, as described in Section 2.1, we find that the observing baseline is not the dominant factor in determining the mass uncertainties at high timing precision. Figures 10 and 11 show how the percent error in the mass-ratio measurements for planets b, c, and d changes for a given simulated survey configuration. For the same simulations (i.e. same realizations of noise), the addition of the third planet to the model greatly reduces the percent errors for planets b and c. For the example case, there is a reduction from 11% to 4% error for the EMB after adding a third planet to our TTVFaster model. Likewise, we can reduce the 20% error on the measurement of Venus’ mass to 8%. In contrast to these two planets, the gas giant’s parameters are usually much less accurate and precise in the 3-planet model (see Fig. 11 and 6).
@@
Figure 12: BICs for $\mathcal{H}_{PP}$ (right panel) and $\mathcal{H}_{PPP}$ (left panel): darker bins have better BIC. For the regions of overlap, the minimum change in BIC is 39, corresponding to a $6.3\sigma$ detection of a Jupiter-analogue. Empty bins indicate where we did not have convergent Markov chains.
@@
Figure 14: Measured planet mass as a function of orbital period for multi-planetary systems with $N_{p}$ planets, listed by discovery method: transit (green circles), RVs (blue squares), TTVs (purple triangles), and others (yellow dots). The mass measurements in this work – for Venus, EMB, Mars, and Jupiter – are taken from the lower panel of Figure 6: the 4-planet model with a total uncertainty of $\approx$ seconds. Based on data from the NASA Exoplanet Archive accessed on 2024 May 29.
@@ 4.1 Model Accuracy
As far as Jupiter is concerned, our results are mixed. In the 3- and 4- planet models, we detect Jupiter at high significance near the correct orbital period for the 30-second, 30-year survey, as well as surveys with shorter durations and larger timing noise levels (see Figure 11). However, the mass of Jupiter is not retrieved accurately in the 3-planet model until high timing precision is achieved (Figure 6), or when a 4-planet model is used (Table 5). This indicates that the detection of a giant planet with TTVs is more straightforward than its characterization. Jupiter’s eccentricity is retrieved more accurately and precisely in the 4-planet model than in the 3-planet model, and its orbital period and transit time (or time of inferior conjunction if it does not transit) are consistent with the correct values in the 4-planet case, but not in the 3-planet case. This indicates that despite its small mass, the presence of Mars does have an effect on the inference of the correct parameters of Jupiter; Mars is small, but it is (relatively) mighty. This can be attributed to the closer proximity of Mars to the Earth-Moon-Barycenter, as well as its proximity to a 2:1 resonance, both of which enhance its dynamical influence on the EMB relative that of Jupiter, whose long orbital period greatly diminishes its dynamical impact on Earth. Nevertheless, Mars is 3.5 orders of magnitude smaller in mass than Jupiter, so after Venus, Jupiter still dominates the TTVs of the EMB, but Mars’ influence does interfere in the 3-planet analysis.
@@ Acknowledgements
Software: CalcEph(Gastineau et al., 2015), TTVFaster (Agol & Deck, 2016), Matplotlib (Hunter, 2007; Caswell et al., 2019), Julia (Bezanson et al., 2017).
@@ A.1 Required Observer Location
$\hat{\textbf{n}}_{\text{obs}}=\pm\frac{\textbf{h}_{2}\times\textbf{h}_{1}}{|\textbf{h}_{2}\times\textbf{h}_{1}|}=\frac{(\textbf{r}_{2}\times\textbf{v}_{2})\times(\textbf{r}_{1}\times\textbf{v}_{1})}{|(\textbf{r}_{2}\times\textbf{v}_{2})\times(\textbf{r}_{1}\times\textbf{v}_{1})|},$ (10)
@@ A.2 Transit Definition
$f\equiv\textbf{r}_{\text{sky,}i}\cdot\textbf{v}_{\text{sky,}i}=0$ (11)
@@
$f(t)=(\textbf{r}_{i}\cdot\textbf{v}_{i})-(\textbf{r}_{i}\cdot\hat{\textbf{n}}_{\text{obs}})(\textbf{v}_{i}\cdot\hat{\textbf{n}}_{\text{obs}}),$ (12)
@@
$f^{\prime}(t)=(\textbf{v}_{i}\cdot\textbf{v}_{i})+(\textbf{r}_{i}\cdot\textbf{a}_{i})-(\textbf{v}_{i}\cdot\hat{\textbf{n}}_{\text{obs}})^{2}\\ -(\textbf{r}_{i}\cdot\hat{\textbf{n}}_{\text{obs}})(\textbf{a}_{i}\cdot\hat{\textbf{n}}_{\text{obs}}).$ (13)

### v2

@@ Abstract
Planetary systems with multiple transiting planets are beneficial for understanding planet occurrence rates and system architectures. Although we have yet to find a solar system (SS) analog, future surveys may detect multiple terrestrial planets transiting a Sun-like star. In this work, we simulate transit timing observations of our system based on the actual orbital motions of Venus and the Earth-Moon Barycenter (EMB) — influenced by the other SS bodies — with varying noise levels and observing durations. We then retrieve the system’s dynamical parameters using an approximate N-body model for transit timing shifts while considering four possible plane-parallel configurations: 2-, 3-, 4-, and 5-planets.
[…]
Astrobiology (74), Exoplanet detection methods (489), Solar system planets (1260)
software: CalcEph(Gastineau et al., 2015), TTVFaster (Agol & Deck, 2016), Matplotlib (Hunter, 2007; Caswell et al., 2019), Julia (Bezanson et al., 2017)
@@ 3 Results
Figure 5: Upper panel: Likelihood profiles for Mars-like (orange curve) and Jupiter-like (firebrick curve) planets fit to the transit times of two terrestrial planets. The times used here were generated over 30 yrs with 30-sec noise. Lower panels: Comparison of likelihood (curve) and posterior probability (histogram) of orbital periods near the measured values for Mars (left) and Jupiter (right). We report the median systematic uncertainty, $\sigma_{sys}$, from the posterior sampling of each configuration, indicating that the four-planet model gives a better fit.
@@ 3.2.4 Five planets
Figure 9: BICs for $\mathcal{H}_{PP}$ (right) and $\mathcal{H}_{PPP}$ (left): darker bins have lower BICs, which indicates a more preferred solution. For the regions of overlap, the minimum change in BIC is 39, corresponding to a $6.3\sigma$ detection of a Jupiter-analogue. Empty bins indicate where we did not have convergent Markov chains.
@@ 3.4 Effect of timing noise and observing baseline
For the $\mathcal{H}_{PPP}$ case with 30-year duration, we plot how timing precision affects our posterior masses in the top panel of Figure 6. Although an increase in injected noise leads to larger mass uncertainties for a given model, the measured masses for planets b and c remain consistent across all of the models. The lower panel shows how timing precision affects the masses in the $\mathcal{H}_{PPPP}$ case, where available.
When we examine our entire simulation set, as described in Section 2.1, we find that the observing baseline is not the dominant factor in determining the mass uncertainties at high timing precision. Figures 11 and 12 show how the percent error in the mass-ratio measurements for planets b, c, and d changes for a given simulated survey configuration. For the same simulations (i.e. same realizations of noise), the addition of the third planet to the model greatly reduces the percent errors for planets b and c. For the example case, there is a reduction from 11% to 4% error for the EMB after adding a third planet to our TTVFaster model. Likewise, we can reduce the 20% error on the measurement of Venus’ mass to 8%. In contrast to these two planets, the gas giant’s parameters are usually much less accurate and precise in the 3-planet model (see Fig. 12 and 6).
@@
Figure 14: Measured planet mass as a function of orbital period for multi-planetary systems with $N_{p}$ planets, listed by discovery method: transit (green circles), RVs (blue squares), TTVs (purple triangles), and others (yellow dots). The mass measurements in this work – for Venus, EMB, Mars, and Jupiter – are taken from the lower panel of Figure 6: the 4-planet model with a total uncertainty of $\approx$ seconds. Based on data from the NASA Exoplanet Archive accessed on 2024 May 29 (NASA Exoplanet Science Institute, 2020).
@@ 4.1 Model Accuracy
As far as Jupiter is concerned, our results are mixed. In the 3- and 4- planet models, we detect Jupiter at high significance near the correct orbital period for the 30-second, 30-year survey, as well as surveys with shorter durations and larger timing noise levels (see Figure 12). However, the mass of Jupiter is not retrieved accurately in the 3-planet model until high timing precision is achieved (Figure 6), or when a 4-planet model is used (Table 5).
This indicates that the detection of a giant planet with TTVs is more straightforward than its characterization. Jupiter’s eccentricity is retrieved more accurately and precisely in the 4-planet model than in the 3-planet model, and its orbital period and transit time (or time of inferior conjunction if it does not transit) are consistent with the correct values in the 4-planet case, but not in the 3-planet case. This indicates that despite its small mass, the presence of Mars does have an effect on the inference of the correct parameters of Jupiter; Mars is small, but it is (relatively) mighty. This can be attributed to the closer proximity of Mars to the Earth-Moon-Barycenter, as well as its proximity to a 2:1 resonance, both of which enhance its dynamical influence on the EMB relative that of Jupiter, whose long orbital period greatly diminishes its dynamical impact on Earth. Nevertheless, Mars is 3.5 orders of magnitude smaller in mass than Jupiter, so after Venus, Jupiter still dominates the TTVs of the EMB, but Mars’ influence does interfere in the 3-planet analysis.
@@ Acknowledgements
This research has made use of the NASA Exoplanet Archive, which is operated by the California Institute of Technology, under contract with the National Aeronautics and Space Administration under the Exoplanet Exploration Program.
@@ A.1 Required Observer Location
$\hat{\textbf{n}}_{\text{obs}}=\pm\frac{\textbf{h}_{2}\times\textbf{h}_{1}}{|\textbf{h}_{2}\times\textbf{h}_{1}|}=\frac{(\textbf{r}_{2}\times\textbf{v}_{2})\times(\textbf{r}_{1}\times\textbf{v}_{1})}{|(\textbf{r}_{2}\times\textbf{v}_{2})\times(\textbf{r}_{1}\times\textbf{v}_{1})|},$ (A1)
@@ A.2 Transit Definition
$f\equiv\textbf{r}_{\text{sky,}i}\cdot\textbf{v}_{\text{sky,}i}=0$ (A2)
@@
$f(t)=(\textbf{r}_{i}\cdot\textbf{v}_{i})-(\textbf{r}_{i}\cdot\hat{\textbf{n}}_{\text{obs}})(\textbf{v}_{i}\cdot\hat{\textbf{n}}_{\text{obs}}),$ (A3)
@@
$f^{\prime}(t)=(\textbf{v}_{i}\cdot\textbf{v}_{i})+(\textbf{r}_{i}\cdot\textbf{a}_{i})-(\textbf{v}_{i}\cdot\hat{\textbf{n}}_{\text{obs}})^{2}-(\textbf{r}_{i}\cdot\hat{\textbf{n}}_{\text{obs}})(\textbf{a}_{i}\cdot\hat{\textbf{n}}_{\text{obs}}).$ (A4)